[Brahms-l] Shift report 20040118 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 18 2004 - 00:00:40 EST
------------------------------------------------------------
Date: Sun Jan 18 00:00:40 2004, Shift: 16:00-24:00
Supervisor: CEJ, with: 
Log book#15 pages 35 - 35

Shift Summary:
------------------------------------------------------------
1) Only one short store from 21:30 to 21:45, with very 
   wide bunches (~35ns). STAR and PHOBOS voted for dump.

2) Pedestal runs: 9733, 9734
   Data run: 9745 (nTr3=16K, nTr6=50K)

3) Still problems with T3 trip. Also T4 and T5 have funny 
   looking summary plots (gaps in the hit dist.).
   H1 slat 16 top tube TDC is dead.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 18 00:01:27 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 18 2004 - 00:01:59 EST